OpenStudy (loser66):

What is \(11^{644}(mod~~645)\) Please, help

OpenStudy (loser66):

Yes, I am here. :)

ganeshie8 (ganeshie8):

645 = 3*5*43

ganeshie8 (ganeshie8):

Notice below : \(11^{644} \equiv (-1)^{644}\equiv 1 \pmod{3}\) \(11^{644} \equiv (1)^{644}\equiv 1 \pmod{5}\) \(11^{644} \equiv (11)^{42*15+14}\equiv 11^{14}\equiv 1 \pmod{43}\)

ganeshie8 (ganeshie8):

that means \(3 \mid (11^{644}-1)\) \(5 \mid (11^{644}-1)\) \(43 \mid (11^{644}-1)\) since \(3, 5, 43\) are primes, it follows \(3*5*43 \mid (11^{644}-1)\) which is same as saying \[11^{644}\equiv 1\pmod{3*5*43}\]

OpenStudy (loser66):

Thank you so much. My question is how we know 11^(14) =1 mod 43 without calculator?

ganeshie8 (ganeshie8):

that is a good question... before that, did you get why \(11^{42*15} \equiv 1\pmod{43}\) ?

OpenStudy (loser66):

Yes

ganeshie8 (ganeshie8):

may i know why you think that reduces to 1 ?

OpenStudy (loser66):

because 43 is a prime, and (43,11) =1 hence \(11^{42}\equiv 1(mod 43)\) and that leads to \(11^ {42*15}\equiv 1(mod 43)\)

OpenStudy (loser66):

Fermat told me that :)

OpenStudy (zarkon):

you shouldn't talk to the dead

OpenStudy (loser66):

He "told" me through the book. :)

ganeshie8 (ganeshie8):

lol nice, do you feel \(11^{14}\) is hard to reduce ?

OpenStudy (loser66):

Yes.

ganeshie8 (ganeshie8):

how ?

ganeshie8 (ganeshie8):

show me how it is hard

ganeshie8 (ganeshie8):

because, i have figured out the answer in 2-3 steps working it in my head...

OpenStudy (loser66):

11^14 = 11*11*11*........(14 times) . I don't know how to make it easier.

ganeshie8 (ganeshie8):

Alright, here it is : \[11^{14} \equiv 121^7 \equiv (-8)^7\equiv - 2^{3*7}\equiv -128^3 \equiv -(-1)^3 \equiv 1\pmod{43}\]

OpenStudy (loser66):

oh, it is much easier. Thank you so much. I got it.